OpenStudy (wilder.monday):

simplify 18x - 6/9x^5 all over 15x + 5/21x^2

OpenStudy (samanthagreer):

ok now cl9ose this and it will just be me and u

OpenStudy (wilder.monday):

ok done

OpenStudy (samanthagreer):

14(3x-1) over 5x^3(3x+1)

OpenStudy (wilder.monday):

explanation again? pls

OpenStudy (samanthagreer):

ok i want you to try 6(3x-1)/9x^5/5(3x+1)/21x^2 now try to simplfy

OpenStudy (wilder.monday):

ok one sec :P

OpenStudy (wilder.monday):

I think what I did Is wrong? I'm so confused ugh. I got up to 6(3x - 1)/3x^5/7x^2/5(3x +1) yeah I have no idea what that is

OpenStudy (samanthagreer):

ok this on i acutaly dont know how to explan it so i will give you a link

OpenStudy (wilder.monday):

ok

OpenStudy (samanthagreer):

hold on i got it we're going to divide (18x - 6)/9x^5 by (15x + 5)/21x^2. the easiest way to do this, is to "flip the 2nd fraction", and multiply: [(18x - 6)/9x^5] * [21x^2/(15x + 5)] now the x^2 on top of the second fraction will cancel partially with the x^5 on the bottom faction, leaving x^3 on the bottom. we can also take a 6 out of 18x - 6 = 6(3x - 1), and a 5 out of 15x + 5 = 5(3x + 1) so: [(18x - 6)/9x^5] * [21x^2/(15x + 5)] = [6(3x - 1)/9x^3] * [21/(5(3x + 1))] = [2(3x - 1)/x^3] * [7/(5(3x + 1)] (3 from the 9 cancelling with 3 from the 6, the other 3 cancelled with the 3 from 21) that's all the cancelling we can do, so now we multiply top times top, and bottom times bottom: = 14(3x - 1)/(5x^3(3x + 1)), which is probably "simple enough" although you can multiply everything together: = (42x - 14)/(15x^4 + 5x^3) sorry i did the math wrong when i did it the first time

OpenStudy (wilder.monday):

ok that makes more sense than what I was doing lol. two more? pls

OpenStudy (samanthagreer):

ok i will message you when i am ready i have to go for a minute ok

OpenStudy (wilder.monday):

alrighty

OpenStudy (samanthagreer):

i promise i will be back